# CLARE: Cognitive Load Assessment in REaltime with Multimodal Data

### v1

@@ Abstract
We present a novel multimodal dataset for Cognitive Load Assessment in REaltime (CLARE). The dataset contains physiological and gaze data from 24 participants with self-reported cognitive load scores as ground-truth labels. The dataset consists of four modalities, namely, Electrocardiography (ECG), Electrodermal Activity (EDA), Electroencephalogram (EEG), and Gaze tracking. To map diverse levels of mental load on participants during experiments, each participant completed four nine-minutes sessions on a computer-based operator performance and mental workload task (the MATB-II software) with varying levels of complexity in one minute segments. During the experiment, participants reported their cognitive load every 10 seconds. For the dataset, we also provide benchmark binary classification results with machine learning and deep learning models on two different evaluation schemes, namely, 10-fold and leave-one-subject-out (LOSO) cross-validation. Benchmark results show that for 10-fold evaluation, the convolutional neural network (CNN) based deep learning model achieves the best classification performance with ECG, EDA, and Gaze. In contrast, for LOSO, the best performance is achieved by the deep learning model with ECG, EDA, and EEG.
@@ I Introduction
For cognitive load assessment, several tasks have been used to introduce varying amounts of mental workload on participants, such as n-back tasks, visual cue tasks, games, arithmetic, multiple-choice tests, and reading exercises, [8, 9, 10]. One software program used for inducing varying levels of cognitive load is the Multi-Attribute Task Battery-II (MATB-II) [11]. Developed by NASA, MATB-II [11] is a computer-based program that can be used as a benchmark to evaluate the workload and performance that aircrews, under conditions similar to flight, might experience [11]. MATB-II has been widely used for assessing emotions and cognitive load in various scientific disciplines [12, 13, 14]. For measuring cognitive load, three general approaches are utilized: self-reporting using tools such as the National Aeronautics and Space Association Task Load Index (NASA TLX) [15] or Subjective Cognitive Load Scores (PAAS) [16], secondary behavioural performance, and the analysis of a range of physiological signals [17].
@@
Mental workload alters sympathetic nervous system activities, which result in changes in Electrocardiogram (ECG), Electrodermal Activity (EDA), also known as Galvanic Skin Response (GSR), gaze, pupil size, Electroencephalogram (EEG), and other physiological responses. These signals can be used to estimate the cognitive load and emotional states using machine learning and deep learning algorithms [18, 19, 20, 21, 22, 23]. Recent advances allow researchers to utilize various machine learning algorithms to classify a user’s cognitive workload using physiological measures. Despite the potential benefits and use cases of automated cognitive load measurement, the lack of annotated cognitive load datasets has been a roadblock towards creating machine learning systems capable of estimating cognitive load with high confidence. Recent studies use self-reported cognitive load as the ground truth for labelling the data; however, cognitive load is generally recorded upon completion of the experiment. This approach can affect the accuracy of the score since it is administered retrospectively, which can introduce inaccuracy due to lapses in memory and recency bias [1]. Also, rather than frequent labels captured during the performed activities, present datasets generally have a single ground truth label associated with long sequences, which can introduce further inaccuracies.
@@
To overcome these problems, we introduce a multimodal dataset on wearable-based Cognitive Load Assessment in REaltime (CLARE). In this dataset, we collected four different data modalities, i.e., ECG, EDA, EEG, and gaze data, from 24 participants while performing MATB-II tasks designed to induce varying amounts of cognitive load. During the experiments, each participant completed four 9-minute sessions, comprised of different cognitive tasks with varying mental effort requirements. We documented the participants’ self-reported subjective cognitive scores on a 9-point Likert scale at 10-second intervals during each session.
As a consequence, our dataset differs from the few cognitive load datasets that have been made publicly available. Our methodology collects subjective self-reported cognitive load scores at frequent intervals throughout the experiments, instead of the commonly used approach of retrospectively recording scores. Given its frequency, this approach better facilitates the training of supervised machine learning solutions, specifically for cognitive load assessment.
@@
    We make the dataset publicly available to facilitate and promote realtime cognitive load estimation research and contribute to the field.
@@
The rest of this paper is organized as follows. In Section II, we give an overview of datasets and work in affective computing and cognitive load analysis by means of biological signals. Section III explains our experimental setup and study protocol. Section IV discusses the data processing procedures and extracted features for classical machine learning algorithms. Sections V and VI present the data analysis and baseline results using the classical machine learning algorithms and deep neural network models. Finally, Section VII concludes the paper.
@@ II-B Cognitive load Datasets
In this section, we review prior work that present datasets on cognitive load assessment. CLAS [33] is a multimodal dataset that collected physiological (ECG, EDA, PPG), and ACC data from 62 participants. During the experiment, to elicit emotions, mental strain, and high cognitive load, the participants performed various tasks such as solving mathematical and logical problems, taking the Stroop test, and watching images and audio-video clips. Gjoreski et al. [8] introduced two multimodal datasets, CogLoad and Snake, for facilitating research on cognitive load inference and personality traits. Both datasets collected physiological data (time interval between R-peaks (RR), EDA, ACC, and ST) from wrist-worn devices from 23 participants. In the CogLoad dataset, participants performed multiple tasks, such as 2- and 3-back tasks. In the Snake dataset, the participants played a game with varying levels of difficulty (easy, medium, and hard) on a smartphone. The participants’ personality traits were recorded using questionnaires in both datasets, and their perceived cognitive load was collected using a self-reporting tool, the National Aeronautics and Space Association Task Load Index (NASA TLX) [15]. Kalatzis et al. [34] introduced a database comprising data from 26 participants for cognitive load research. ECG and RR data were collected from the participants while they performed two tasks to induce low and high cognitive workloads using the MATB-II software. The participants’ perceived cognitive load levels were recorded using the NASA-TLX.
@@ III-A MATB-II Software
To induce varying amounts of cognitive load, we used the Multi-Attribute Task Battery (MATB-II) tool [11] developed by the National Aeronautics and Space Administration (NASA) [^1] https://www.nasa.gov/ [Accessed: 2022-06- 05]. This tool was developed to facilitate research on human multi-task performance and mental workload. As shown in Figure 1, the MATB-II interface has five regions comprising four tasks and one scheduling region. These regions are explained as follows:
@@
    Tracking task: In this task, shown in region 2 of Figure 1, there are two modes: automatic and manual. In the automatic mode, no attention is required from the participant, while in the manual mode, the participant needs to actively use the joystick to keep the crosshair inside the blue square in the middle.
@@
    Resource management task: In this task, shown in region 4 in Figure 1, the goal is to keep tanks A and B at the specified levels (marked with blue bars on the side of tanks). These tanks keep depleting during the experiment and must be refilled using tanks C, D, E, and F. The participant can circulate the fuel in the tanks using pumps marked as 1-8. The arrow ‘$>$’ shows the direction in which the pumps can circulate fuel. Tanks C and D have limited capacity but tanks E and F have unlimited capacity. The pumps can be toggled to ‘inactive’ (idle shown with white colour) or ‘active’ (pumping shown with green) states by clicking on them. The pumps can also break (red colour) during the experiment, making the pump unusable. After a specific time, the broken pumps are fixed automatically.
@@
Figure 2: Sensor placements, ECG (in blue), EDA (in red), and EEG (in green).
@@ III-C Sensors and Placement
Finally, Tobii Pro Glasses 2777https://www.tobiipro.com/product-listing/tobii-pro-glasses-2/ [Accessed: 2022-05-21], shown in Figure 5(d), were used to record the direction and velocity of the gaze. The glasses consist of two units, a head unit and a recording unit. The gaze-tracking data was recorded at a sampling rate of 50 Hz.
@@ III-D Experiment Protocol
After the participant was seated, the wireless connection to the sensors was established using a remote computer (Dell Alienware 15 laptop) for data capture via Bluetooth. An additional Bluetooth adapter was used for better connectivity and to reduce data packet loss. Then a member of the research team provided the participant with a guided walkthrough of the experiment, including the purpose of research, sequence of events, expected time commitments, and any procedural tips that could help the participant navigate the process. The layout of the experiment area is shown in Figure 6.
@@
As mentioned above, each participant was asked to complete four 9-minute MATB-II sessions, the first of which was the linearly ascending complexity session. Participants provided their subjective cognitive load scores verbally during each session by responding to a beep sound after every 10-seconds. After each 9-minute session, each participant was given 2-minutes of resting time. Figure 7 shows a participant during the experiment. At the end of the fourth session, participants completed a debriefing discussion by answering some qualitative questions about the experiment. All participants received a monetary compensation of 70 CAD after the experiment.
@@
Figure 7: (a) Experiment setup front view, and (b) experiment setup side view
@@ IV-A Data Preprocessing
ECG Time Domain Min HR, Min HRV, Min ECG, Max HR, Max HRV, Max ECG, Mean HR, Mean HRV, Mean ECG, SD HR, SD HRV, SD, ECG, RMS HRV, % HRV $>$ 50 ms, % HRV $>$ 20 ms, IQR HRV, IQR ECG, MAD HRV, MAD ECG, Skewness ECG, kurtosis ECG, entropy ECG, AUC2 ECG
@@
Block 1 Conv1D, 64, 32, $\begin{array}[]{cc}\text{Conv1D, 64, 32, 1}\\ \text{Conv1D, 64, 32, 3}\\ \end{array}$
Block 2 Conv1D, 32, 64, $\begin{array}[]{cc}\text{Conv1D, 32, 64, 1}\\ \text{Conv1D, 32, 64, 3}\\ \end{array}$
Block 3 Conv1D, 17, 128, $\begin{array}[]{cc}\text{Conv1D, 17, 128, 1}\\ \text{Conv1D, 17, 128, 3}\\ \end{array}$
Block 4 Conv1D, 7, 256, $\begin{array}[]{cc}\text{Conv1D, 7, 256, 1}\\ \text{Conv1D, 7, 256, 3}\\ \end{array}$
@@ IV-B Feature Extraction
ECG Features: We extracted various statistical, time domain and frequency domain features from 10-second segments of filtered ECG signals. The features included time domain features such as mean, standard deviation, maximum and minimum, and median of absolute values of HR, HRV, and filtered ECG. Additionally, skewness, kurtosis, entropy, and AUC2, IQR, pNN20, and pNN50 were extracted. Frequency domain features were also extracted using Welch’s method for power spectral analysis and included peak frequencies, absolute powers, normalized powers, and the LF/HF ratio of ultra-low frequency (ULF) band ($<0.003Hz$), very low frequency (VLF) band (0.003 Hz - 0.04 Hz), low frequency (LF) band (0.04 Hz - 0.15 Hz), and high frequency (HF) band (0.15 Hz - 0.4 Hz), as well as the total power over all frequency bands.
@@ IV-C Dataset Release
To obtain access to the dataset, please contact prithila.angkan@queensu.ca
@@ V Data Analysis
Figure 9: Complexity vs cognitive load scores. The darker shades mean more samples.
@@ VI-A Classification Algorithms
We used eight classical machine learning algorithms and a deep learning model for benchmarking. The features extracted from the four modalities in subsection IV-B served as inputs to the machine learning algorithms for classifying cognitive load scores. The eight classical machine learning algorithms are Gradient Boosting (GB), Light Gradient Boosting Machine (LGBM), Linear Discriminant Analysis (LDA), Logistic Regression (LR), Multilayer Perceptron (MLP), Random Forest (RF), Support Vector Machine (SVM), and Extreme Gradient Boosting (XGBoost). The deep learning benchmark is a VGG-style convolutional neural network (CNN).
@@
For the GB classifier, the number of base estimators was set to 300, the loss is set to logarithmic loss, and a maximum depth of 3 was used for individual estimators. For the LGBM, the number of estimators was set to 2000, and for each base estimator, the number of leaves was set to 100. A learning rate of 0.001 was used. The ‘least squares solution‘ was selected based on hyperparameter tuning for the LDA classifier. For the LR classifier, the maximum number of iterations required for the solvers to converge was set to 400, and the inverse regularization parameter ‘C’ was set to 1. The MLP classifier was trained with two hidden layers of sizes 100 and 10, and an adaptive learning rate for 1000 iterations. The number of estimators for the RF classifier was set to 1000 with a minimum sample split of 5 and a maximum depth of 5 for each tree. The SVM classifier was trained with the inverse regularization parameter ‘C’ set to 10. For the XGBoost classifier, we used a tree-based boosting method with 300 estimators. We set the learning rate and the L1 regularization values to 1e-3 and 1e-4, respectively. The models’ parameters were tuned empirically to obtain the best training results.
@@
VGG-like encoder blocks were used for the deep learning model, which contained two 1D convolutional layers with a ReLU activation layer and a MaxPool layer of filter size and stride of 2. The output of the fourth encoder block was fed to two FC layers with 512 and 256 units, respectively. A feature-level fusion strategy (late fusion) was then used to achieve a multimodal setup. The final output class was generated by SoftMax following an FC layer. Figure 10 shows the multimodal pipeline with the four recorded modalities. The details of filters, filter sizes, and strides for each encoder block are presented in Table III. Similar to the machine learning models, the hyperparameters of the CNN were empirically tuned to achieve the best performance.

### v2

# CLARE: Cognitive Load Assessment in Real-time with Multimodal Data

@@ Abstract
We present a novel multimodal dataset for Cognitive Load Assessment in REal-time (CLARE). The dataset contains physiological and gaze data from 24 participants with self-reported cognitive load scores as ground-truth labels. The dataset consists of four modalities, namely, Electrocardiography (ECG), Electrodermal Activity (EDA), Electroencephalogram (EEG), and Gaze tracking. To map diverse levels of mental load on participants during experiments, each participant completed four nine-minutes sessions on a computer-based operator performance and mental workload task (the MATB-II software) with varying levels of complexity in one minute segments. During the experiment, participants reported their cognitive load every 10 seconds. For the dataset, we also provide benchmark binary classification results with machine learning and deep learning models on two different evaluation schemes, namely, 10-fold and leave-one-subject-out (LOSO) cross-validation. Benchmark results show that for 10-fold evaluation, the convolutional neural network (CNN) based deep learning model achieves the best classification performance with ECG, EDA, and Gaze. In contrast, for LOSO, the best performance is achieved by the deep learning model with ECG, EDA, and EEG.
@@ I Introduction
For cognitive load assessment, several tasks have been used to introduce varying amounts of mental workload on participants, such as n-back tasks, visual cue tasks, games, arithmetic, multiple-choice tests, and reading exercises, [8, 9, 10]. For measuring cognitive load, three general approaches are utilized: self-reporting using tools such as the National Aeronautics and Space Association Task Load Index (NASA TLX) [11] or Subjective Cognitive Load Scores (PAAS) [12], secondary behavioural performance, and the analysis of a range of physiological signals [13].
Prior work has shown that cognitive load and affect are closely related and can influence each other [14]. One recent study showed the transferability of learned representations between these two domains [15], where emotion datasets were used for pre-training encoders which were then used for downstream cognitive load classification. This study indicated the close relationship between cognitive load and affect from a computational perspective. Similarly, from a physiological perspective, the link between cognitive load and affect has been well-established [16, 17, 18, 19]. Mental workload alters the sympathetic nervous system activities, which results in changes in Electrocardiogram (ECG), Electrodermal Activity (EDA), also known as Galvanic Skin Response (GSR), gaze, pupil size, Electroencephalogram (EEG), and other physiological responses. These signals can be used to estimate the cognitive load and emotional states using machine learning and deep learning algorithms [20, 21, 22, 23, 24, 25] and can be easily monitored using off-the-shelf wearables.
@@
Recent advances allow researchers to utilize various machine learning algorithms to classify a user’s cognitive workload using physiological measures. Despite the potential benefits and use cases of automated cognitive load measurement, the lack of annotated cognitive load datasets has been a roadblock towards creating machine learning systems capable of estimating cognitive load with high confidence. Recent studies use self-reported cognitive load scores as ground truth labels; however, these scores are generally recorded upon completion of the experiment, and are thus unable to capture changes in cognitive load throughout the data collection process. This approach can affect the fidelity of the labels since it is administered retrospectively, which can introduce inaccuracies due to lapses in memory and recency bias [1]. Additionally, the lack of frequent labels captured during the performed activities inhibits the development of machine learning systems capable of real-time estimation of cognitive load.
@@
To overcome these problems, we introduce a multimodal dataset on wearable-based Cognitive Load Assessment in REaltime (CLARE). In this dataset, we collected four different data modalities, i.e., ECG, EDA, EEG, and gaze data, from 24 participants while performing MATB-II tasks designed to induce varying amounts of cognitive load. During the experiments, each participant completed four 9-minute sessions, comprised of different cognitive tasks with varying mental effort requirements. We documented the participants’ self-reported subjective cognitive scores on a 9-point Likert scale at 10-second intervals during each session. As a consequence, our dataset differs from the few cognitive load datasets that have been made publicly available. Our methodology collects subjective self-reported cognitive load scores at frequent intervals throughout the experiments, instead of the commonly used approach of retrospectively recording scores. Given its frequency, this approach better facilitates the development of real-time machine learning solutions for cognitive load assessment.
@@
    We make the dataset publicly available to facilitate and promote real-time cognitive load estimation research and contribute to the field.
@@
CL-Drive [37] 2024 21 Cognitive load ECG, EDA, EEG, Gaze Driving simulator
@@ II-B Cognitive Load Datasets
In this section, we review prior work that present datasets on cognitive load assessment. CLAS [35] is a multimodal dataset that collected physiological (ECG, EDA, PPG), and ACC data from 62 participants. During the experiment, to elicit emotions, mental strain, and high cognitive load, the participants performed various tasks such as solving mathematical and logical problems, taking the Stroop test, and watching images and audio-video clips. Gjoreski et al. [8] introduced two multimodal datasets, CogLoad and Snake, for facilitating research on cognitive load inference and personality traits. Both datasets collected physiological data (time interval between R-peaks (RR), EDA, ACC, and ST) from wrist-worn devices from 23 participants. In the CogLoad dataset, participants performed multiple tasks, such as 2- and 3-back tasks. In the Snake dataset, the participants played a game with varying levels of difficulty (easy, medium, and hard) on a smartphone. The participants’ personality traits were recorded using questionnaires in both datasets, and their perceived cognitive load was collected using a self-reporting tool, the National Aeronautics and Space Association Task Load Index (NASA TLX) [11]. Kalatzis et al. [36] introduced a database comprising data from 26 participants for cognitive load research. ECG and RR data were collected from the participants while they performed two tasks to induce low and high cognitive workloads using the MATB-II software. The participants’ perceived cognitive load levels were recorded using the NASA-TLX. CL-Drive [37], used four modalities EEG, ECG, EDA and Gaze data to classify cognitive load during driving in a vehicle simulator. The data was collected from 21 participants and the self-reported cognitive load scores were recorded every 10 seconds. The dataset contains 9 trials of 3 minuted each for each participant of varying cognitive load.
Overall, our review of the existing datasets on cognitive load reveals a number of limitations, which we address through CLARE in this paper. Specifically, we observe that some datasets such as [36] include a low number of modalities, whereas our dataset offers large number and diverse modalities for cognitive load classification. Additionally, most datasets in this area, for instance [36, 35, 8], record cognitive load scores at the end of the experiment, and do not provide labels that correspond the wearable signals throughout the experiments. This means that most existing datasets cannot be used to develop machine learning systems capable of real-time classification of cognitive load. Our dataset, on the other hand, contains frequent ground-truth cognitive load scores recorded throughout the data collection process, which could enable the development of real-time systems.
@@ III-A MATB-II Software
To induce varying amounts of cognitive load, we used the Multi-Attribute Task Battery (MATB-II) tool [39] developed by the National Aeronautics and Space Administration (NASA) [^1] https://www.nasa.gov/ [Accessed: 2022-06-05]. MatB-II is a widely used software by researchers to induce cognitive load in participants. This tool was developed to facilitate research on human multi-task performance and mental workload and has been widely used for assessing emotions and cognitive load in various scientific disciplines [40, 41, 42, 43, 44]. MATB-II is able to induce different levels of cognitive load in participants by changing the complexity of the tasks performed. It consists of 5 different tasks which can be active individually or at the same time to control the complexity, making it ideal for our application. As shown in Figure 1, the MATB-II interface has five regions comprising four tasks and one scheduling region. These regions are explained as follows:
@@
    Tracking task: In this task, shown in region 2 of Figure 1, there are two modes: automatic and manual. In the automatic mode, no attention is required from the participant, while in the manual mode, the participant needs to actively use the joystick to keep the crosshair inside the black square in the middle.
@@
    Resource management task: In this task, shown in region 4 in Figure 1, the goal is to keep tanks A and B at the specified levels (marked with black bars on the side of tanks). These tanks keep depleting during the experiment and must be refilled using tanks C, D, E, and F. The participant can circulate the fuel in the tanks using pumps marked as 1-8. The arrow ‘>>’ shows the direction in which the pumps can circulate fuel. Tanks C and D have limited capacity but tanks E and F have unlimited capacity. The pumps can be toggled to ‘inactive’ (idle shown with white colour) or ‘active’ (pumping shown with green) states by clicking on them. The pumps can also break (red colour) during the experiment, making the pump unusable. After a specific time, the broken pumps are fixed automatically.
@@
Figure 2: Sensor placements, ECG (in black), EDA (in red), and EEG (in green).
@@ III-C Sensors and Placement
Finally, Tobii Pro Glasses 2777https://www.tobiipro.com/product-listing/tobii-pro-glasses-2/ [Accessed: 2022-05-21], shown in Figure 5(d), were used to record the direction and velocity of the gaze. The glasses consist of two units, a head unit and a recording unit. The gaze-tracking data was recorded at a sampling rate of 50 Hz. All the data has been collected and synchronized using the iMotion [^8] https://imotions.com/ software
@@ III-D Experiment Protocol
After the participant was seated, the wireless connection to the sensors was established using a remote computer (Dell Alienware 15 laptop) for data capture via blacktooth. An additional blacktooth adapter was used for better connectivity and to reduce data packet loss. Then a member of the research team provided the participant with a guided walkthrough of the experiment, including the purpose of research, sequence of events, expected time commitments, and any procedural tips that could help the participant navigate the process. The layout of the experiment area is shown in Figure 6.
@@
As mentioned above, each participant was asked to complete four 9-minute MATB-II sessions, the first of which was the linearly ascending complexity session. Participants provided their subjective cognitive load scores verbally during each session by responding to a beep sound after every 10-seconds. After each 9-minute session, each participant was given 2-minutes of resting time. At the end of the fourth session, participants completed a debriefing discussion by answering some qualitative questions about the experiment. All participants received a monetary compensation of 70 CAD after the experiment.
@@ IV-A Data Preprocessing
ECG Time Domain Min HR, Min HRV, Min ECG, Max HR, Max HRV, Max ECG, Mean HR, Mean HRV, Mean ECG, SD HR, SD HRV, SD, ECG, RMS HRV, % HRV >> 50 ms, % HRV >> 20 ms, IQR HRV, IQR ECG, MAD HRV, MAD ECG, Skewness ECG, kurtosis ECG, entropy ECG, AUC2 ECG
@@
[…]
@@ IV-B Feature Extraction
ECG Features: We extracted various statistical, time domain and frequency domain features from 10-second segments of filtered ECG signals. The features included time domain features such as mean, standard deviation, maximum and minimum, and median of absolute values of HR, HRV, and filtered ECG. Additionally, skewness, kurtosis, entropy, and AUC2, IQR, pNN20, and pNN50 were extracted. Frequency domain features were also extracted using Welch’s method for power spectral analysis and included peak frequencies, absolute powers, normalized powers, and the LF/HF ratio of ultra-low frequency (ULF) band (<0.003Hz<0.003Hz), very low frequency (VLF) band (0.003 Hz - 0.04 Hz), low frequency (LF) band (0.04 Hz - 0.15 Hz), and high frequency (HF) band (0.15 Hz - 0.4 Hz), as well as the total power over all frequency bands.
@@ IV-C Dataset Release
We make the dataset public at:
@@ V Data Analysis
Figure 8: Complexity vs. cognitive load scores. The darker shades mean more samples.
@@ VI-A Classification Algorithms
We use eight classical machine learning algorithms and two deep learning models for benchmarking. The features extracted from the four modalities (described in Section IV-B) serve as inputs to all the models. The eight classical machine learning algorithms are Gradient Boosting (GB), Light Gradient Boosting Machine (LGBM), Linear Discriminant Analysis (LDA), Logistic Regression (LR), Multilayer Perceptron (MLP), Random Forest (RF), Support Vector Machine (SVM), and Extreme Gradient Boosting (XGBoost). The deep learning models are a VGG-style convolutional neural network (CNN) and a Transformer network. The selected classical machine learning models have shown strong performances in the past for affective computing from time-series [54, 55], and each contain different strengths and weaknesses [56]. Similarly, the CNN and Transformer models used are popular deep learning techniques that have been successfully used for affective computing with wearable signals in the past [57, 58, 59], making them reliable choices for benchmarking our dataset’s performance. In particular, the selected architectures for the CNN and Transformer have been widely used in prior works [57, 58, 59].
@@
For the GB classifier, the number of base estimators is set to 300, the loss is set to logarithmic loss, and a maximum depth of 3 is used for individual estimators. For the LGBM, the number of estimators is set to 2000, and for each base estimator, the number of leaves was set to 100. A learning rate of 0.001 is used. The ‘least squares solution‘ is selected based on hyperparameter tuning for the LDA classifier. For the LR classifier, the maximum number of iterations required for the solvers to converge is set to 400, and the inverse regularization parameter ‘C’ is set to 1. The MLP classifier is trained with two hidden layers of sizes 100 and 10, and an adaptive learning rate for 1000 iterations. The number of estimators for the RF classifier was set to 1000 with a minimum sample split of 5 and a maximum depth of 5 for each tree. The SVM classifier was trained with the inverse regularization parameter ‘C’ set to 10. For the XGBoost classifier, we used a tree-based boosting method with 300 estimators. We set the learning rate and the L1 regularization values to 1e-3 and 1e-4, respectively. The models’ parameters were tuned empirically to obtain the best training results.
@@
The CNN consists of 4 convolution blocks, each containing two 1D convolutional layers with a ReLU activation layer and a MaxPool layer of filter size and stride of 2. The output of the fourth encoder block is fed to two FC layers with 512 and 256 units, respectively. A feature-level fusion strategy (late fusion) is then used to create the multimodal setup. The final output class is generated by SoftMax following an FC layer. Figure 9 shows the multimodal pipeline with the four recorded modalities. The details of filters, filter sizes, and strides for each encoder block are presented in Table III. Similar to the machine learning models, the hyperparameters of the CNN were empirically tuned to achieve the best performance.
The Transformer network [60] consists of four Transformer blocks and a prediction head. The output from the Transformer blocks is flattened and fed into the prediction head. The prediction head contains two prediction blocks consisting of FC layers of size 256 and 128 followed by ReLU and dropout of 0.5. The output layer contains a sigmoid function. For the multimodal setup, all the features from the individual modalities are concatenated before feeding the Transformer network (early fusion). We used an early fusion setup here as it yielded better results in comparison to the late fusion setup used for the CNN. The entire sequence is then considered as one patch in order to capture the global context of the entire sequence and processed all together by the Transformer network.
@@
Figure 10: Multimodal pipeline of the transformer network.
VI-B Evaluation Protocol and Implementation Details
For all the benchmarks, we use binary (‘high’ versus ‘low’) cognitive load classification. To calculate binary labels, reported scores of less than 5 were considered ‘low’ cognitive load, and scores of 5 or more were considered ‘high’ cognitive load. This threshold is chosen because it provides a clear and straightforward division between lower and higher levels of cognitive load, allowing us to perform binary classification. Our post-collection interviews with participants indicated that many participants attributed 5 with a relatively high cognitive load state, and therefore we placed 5 and higher in the ‘high’ category, while 1 to 4 were placed in the ‘low’ category. We use 10-fold and Leave-One-Subject-Out (LOSO) cross-validation evaluation schemes to indicate the performance of our models on previously unseen data. Accuracy and F1 score were used as evaluation metrics. For all the machine learning algorithms except LGBM and XGBoost, we use the Scikit-learn [^9] https://scikit-learn.org/stable/ library. All the parameters of the machine learning algorithms were optimized by performing an exhaustive search. Our deep learning pipelines were implemented using the TensorFlow framework. For the CNN, we used AdaDelta with a decay rate of 0.95, a learning rate of 5e-3, and focal loss (alpha of 4.0 and gamma of 2.0). For the transformer network we used Adam [61] optimizer with a learning rate of 0.0001. We empirically selected a batch size of 256 for 100 epochs for all experiments. We implemented our pipelines using an Intel Core i7-9700 CPU and an NVIDIA GeForce RTX 2080 Ti GPU.